The length of a rectangle is three less than twice the width. If the area is 54 what is the length and width of the rectangle?
Answer by mananth(16946)

width = w
length = 2w-3
..
Area = L*W
w(2w-3)=54
2w^2-3w=54
2w^2-3w-54=0
2w^2-12w+9w-54=0
2w(w-6)+9(w-6)=0
(w-6)(2w+9)=0
w= 6 the width
length = 2w-3
=12-3
length = 9
...
m.ananth@hotmail.ca